coffe worth 90.00 per klo. is mix with a coffee worth 75.00 per klo. to produce 300 klo. grams of the mixture worth 80.00 per klo. how many klos. of each type should she use?
Answer by ptaylor(2198)

Let x=amount of 90/kilo needed
Then (300-x)=amount of 75/kilo coffee needed
Now we know that the value of 90/kilo coffee (90x) plus the value of the 75/kilo coffee (75(300-x)) has to equal the value of the final mixture(80)(300)). So our equation to solve is:
90x+75(300-x)=80*300 get rid of parens
90x+22500-75x=24000 subtract 22500 from both sides
90x-75x+22500-22500=24000-22500 collect like terms
15x=1500 divide both sides by 15
x=100 kilos-------------------------amount of 90/kilo coffee needed
300-x=300-100=200 kilos---------------amount of 75/kilo coffee needed
CK
90*100+75*200=80*300
9000+15000=24000
24000=24000
Hope this helps---ptaylor